A stock has a beta of 1.80 and an expected return of 13 percent. A risk-free asset currently earns 3.2 percent.

a. What is the expected return on a portfolio that is equally invested in the two assets?

b. If a portfolio of the two assets has a beta of .99, what are the portfolio weights?

c. If a portfolio of the two assets has an expected return of 9%, what is its beta?

d. If a portfolio of the two assets has a beta of 3.60, what are the portfolio weights?
Answer #1

a.

Expected Rate = 0.50(0.13) + 0.50(0.032)

Expected Rate = 8.10%

b

0.99 = 1.80w

w = 0.55

55% in stock and 45% in risk free asset

c.

Using CAPM Model,

Rm = 0.032 + (0.13 - 0.032)/1.8 = 8.64%

Beta = (0.09 - 0.032)/(0.0864 - 0.032)

Beta = 1.07

d.

3.60 = 1.80w

w = 2

Weight of Stock = 200%

Weight of Risk Free Asset = -100%